OpenStudy (cherry17ann):

if a piece of clay has the measurements 2.0 cm x 4.0 x 6.0 x cm and weighs 96.00 grams, calculate its density

OpenStudy (otherworldly):

denisity = mass divided by volume

OpenStudy (cherry17ann):

Is the mass all the numbers calculated except grams?

OpenStudy (otherworldly):

are all those numbers length width and height

OpenStudy (otherworldly):

except the grams?

OpenStudy (otherworldly):

@cherry17ann

OpenStudy (otherworldly):

if they r then they r volume so 96/48 so 2g/cm

OpenStudy (cherry17ann):

CN you explain how you got that please?

OpenStudy (cherry17ann):

Oh nevermind ... i think i got it thank you so much

OpenStudy (otherworldly):

np

OpenStudy (danjs):

Density is a mass per unit of volume of something.. \[\rho = \frac{ m }{ V }\] \[\rho = \frac{ 96.0~~g }{ 2.0~cm * 4.0~cm *6.0~cm }=\frac{ 96.0~~g }{ 48.0~~cm^3 }=\frac{ 2.0~~~g }{ cm^3 }\]